# Enrich the content of the image Using Context-Aware Copy Paste

### v1

@@ Introduction
To evaluate the performance of the object detection model, we use mAP as out metric, which stands for Mean Average Precision. Here we set the threshold as 50%, which means IoU over 0.5 will be considered as
@@
Evaluation Datasets
To comprehensively assess the viability of our proposed approach, we conducted experiments across three distinct datasets: Cat-Dog classification(?), and CityPersons(?). These datasets are representative of key tasks in the computer vision field: classification, segmentation, and object detection, respectively.
The Cat-Dog dataset contains 25,000 images, each labeled as either a cat or a dog. The CityPersons(?) dataset is a subset of Cityscapes, focusing solely on person annotations. It includes 2,975 images for training and 500 images for validation. The Cambridge driving Labeled Video Database(CamVid)(?) is the first collection of videos with object class semantic labels, complete with metadata. The dataset contain over 700 images with pixel-level annotation. The annotation of images cover 32 class labels, include building, car, tree, sky, etc.
@@
Results
Across initialization
@@ Results
To valid the robustness of CACP across different initialization, we conduct experiments on CamVid based on two different initialization configurations, namely imagenet(?) pretrained and normal initialization. As illustrated in Fig .5, the results with CACP outperform the one without CACP in both configurations: 0.803 to 0.817 in random parameter initialization; 0.842 to 0.864 in imagenet pre-trained config.
Figure 5: Comparison between different initialization configuration W/Wo CACP.
Across Tasks
@@
We conduct experiments on different computer vision tasks to valid the usage scenarios of CACP, namely image classification tasks on Cat-Dog, image segmentation and object detection on CityPersons. The item in column Methods indicate different combinations. B indicates base model. In classification task, which is Resnet-50, in segmentation is DeepLabv3, and in object detection is YOLOv5-s. CP and CACP indicate random crop-paste and context-aware copy-paste, respectively. aug is a combination of traditional data augmentation techniques, including flip, colorjitting, random noise, which is provided by lib Albumentation. As illustrated in Table 5, all augmentation techniques have the ability to boost the model. CACP outperforms CP and aug across all three tasks.
Across Architectures
@@
To valid the effectiveness of our method across different architectures, we conduct experiments on CamVid(?) dataset using popular encoder-decoder architectures, namely U-net(?), FPN(?), PSPNet(?), DeepLabv3(?), DeepLabV3+(?), and PAN(?). The experiment is conducted with or without CACP operation. To observe the effect of different categories, we select six classes: [car,pedestrian,building,tree,sky,road].
Across Partition
@@
To valid the effect of the number of augmented images, we conduct the experiments on CamVid. 1/n indicates 1/n of training images have been augmented using CACP. It is important to note that the total number of training image is fixed, only the ration of augmented and non-augmented varies. The result illustrates that the performance increases with the rise of partition from 1/8 to 1/2, the trend is stable in all 4 experiments. The increase is saturated when partition is over 1/2.
Speed up convergence
@@
We have noticed that CACP contributes to speed up the training process. As illustrated in Fig .6, the CACP augmented one converges rapidly compared to the wo-CACP one. The loss is stable around epoch 15, while the wo-CACP is still not fully converged after epoch 19.
@@
Figure 6: dice loss during training with or without CACP.
Effect of GradCAM Prompt
@@
To improve the robustness and stability of the output masks of SAM, we propose the GradCAM-guided approach to generate points+bounding box prompts. During our trials, we notice that how many points in GradCAM to select affect the final segmentation mask. To determine the best point numbers, we conduct the following experiment on CamVid. random indicates the point is sampled randomly inside bounding box. CAM(n) means extract n points in GradCAM map with high value. The Table7 indicates that extra prompts can improve the accuracy of the SAM; CAM-based point prompts are better than random points. The prefered number of points is around 3 to 5.
@@
Table 7: mIoU across different prompts.
@@
Discussion
n this paper, we propose a context-aware copy-paste (CACP) data augmentation approach that serves as a versatile plug-in module for various computer vision tasks, eliminating the need for additional manual annotation. CACP proves highly efficient for custom segmentation tasks in specific scenarios. Instead of annotating thousands of pixel-level masks, fewer than 100 annotated base images are sufficient to generate a multitude of highly realistic pseudo pixel-level images for training models. This process is both time-saving and scalable, allowing users to customize their target gallery according to specific task requirements.
As a case study, we applied CACP to our Railway obstacle detection project (?). With just a single annotation mask (due to fixed camera placement), we trained a universal obstacle detection model that outperformed previous object detection-based approaches. CACP proves particularly effective in semi-supervised segmentation by generating high-quality pseudo images. Moving forward, we aim to develop more user-friendly tools to enhance accessibility and extendability.

### v2

@@ 5.2 Metric
To evaluate the performance of the object detection model, we use (mean Average Precision)mAP as our metric. Here we set the threshold as 50%, which means IoU over 0.5 will be considered as correct detection.
@@
5.3 Evaluation Datasets
To comprehensively assess the viability of our proposed approach, we conducted experiments across two distinct datasets: Cat-Dog classification[25], and CityPersons[31]. These datasets are representative of key tasks in the computer vision field: classification, segmentation, and object detection, respectively.
The Cat-Dog dataset contains 25,000 images, each labeled as either a cat or a dog. The CityPersons[31] dataset is a subset of Cityscapes, focusing solely on person annotations. It includes 2,975 images for training and 500 images for validation. The Cambridge driving Labeled Video Database(CamVid)[2] is the first collection of videos with object class semantic labels, complete with metadata. The dataset contains over 700 images with pixel-level annotation. The annotation of images cover 32 class labels from urban and non-urban driving scenes.
@@ 5.3 Evaluation Datasets
5.4 Results
5.4.1 Across Initialization
@@
To validate the robustness of CACP across different initialization, we conduct experiments on CamVid based on two different initialization configurations, namely ImageNet[9] pretrained and normal initialization. As illustrated in Fig .5, the results with CACP outperform the one without CACP in both configurations.
Figure 5: CACP provides gains that are robust to training configurations. We train DeepLabv3 on CamVid for varying number of epochs. The CACP is helpful under with and without pretraining configurations.
5.4.2 Across Tasks
@@
We conduct experiments on different computer vision tasks to validate the usage scenarios of CACP, namely image classification tasks on Cat-Dog, image segmentation and object detection on CityPersons. As illustrated in Table 4, the items in the column Methods indicate different combinations. B indicates base model. In the classification task this is a Resnet-50 and in segmentation it is DeepLabv3[4], while in object detection it is YOLOv5-s[30]. CP and CACP indicate random crop-paste and context-aware copy-paste, respectively. aug is a combination of traditional data augmentation techniques, including flip, Color jittering, random noise, which is provided by lib Albumentation[3]. All augmentation techniques have the ability to boost the model. CACP outperforms CP and aug across all three tasks.
5.4.3 Across Architectures
@@
As illustrated in Table 3, to validate the effectiveness of our method across different architectures, we conduct experiments on CamVid[1] dataset using popular encoder-decoder architectures, namely U-net[27], FPN[19], PSPNet[32], DeepLabv3[4], DeepLabV3+[5], and PAN[21]. The experiment is conducted with or without our CACP augmentation operation. To observe the effect of different categories, we select six classes: [car,pedestrian,building,tree,sky,road]. It can be noticed that almost all classes results are improved with CACP augmentation (except Sky in DeepLabV3 configuration).
5.4.4 Across Partition
@@
To validate the effect of the number of augmented images, we conduct experiments on CamVid as illustrated in Table 5. 1/n indicates 1/n of training images have been augmented using CACP. It is important to note that the total number of training image is fixed, only the ratio of augmented to non-augmented varies. The result illustrates that the performance increases with the rise of partition from 1/8 to 1/2, the trend is stable in all 4 experiments. The increase is saturated when partition is over 1/2.
5.4.5 Speed Up Convergence
@@
We have noticed that CACP contributes to speed up the training process. We trained DeepLabv3 on Camvid with 20 epochs. As illustrated in Fig. 6, the CACP augmented one converges rapidly compared to the wo-CACP one. The loss is stable around epoch 15, while the wo-CACP is still not fully converged after epoch 19.
@@
Figure 6: During training, the Dice loss with the CACP configuration converges faster compared to the process without the CACP configuration.
5.4.6 Effect of GradCAM Prompt
@@
To improve the robustness and stability of the output masks of SAM, we propose the GradCAM-guided approach to generate points+bounding box prompts. During our trials, we notice that the number of selected points in GradCAM affects the final segmentation mask. To determine the best point numbers, we conduct the following experiment on CamVid. Rand indicates the point is sampled randomly inside bounding box. CAM(n) means n points are extracted in GradCAM map with high value. Table 6 indicates that extra prompts can improve the accuracy of SAM; CAM-based point prompts are better than random points. The preferred number of points is around 3 to 5.
@@
Table 6: mIoU across different prompts.The performance improves with the number of CAM-generated point prompts and stabilizes when the number of point prompts exceeds three.
@@
6 Discussion
In this paper, we propose a context-aware copy-paste (CACP) data augmentation approach, designed as a versatile plug-and-play module for various computer vision tasks, eliminating the need for additional manual annotation. CACP is particularly effective for custom segmentation in semi-supervised learning, offering a time-efficient and scalable solution that allows users to tailor their target gallery to specific task requirements.
Future research directions include the following:
  • Integration with diffusion models to generate synthetic augmented images with well-crafted prompts, particularly for privacy-sensitive objects, thereby mitigating privacy concerns.
  • Adaptation to downstream industrial applications, such as obstacle detection and pedestrian detection.
  • Enhancement through advanced image composition techniques, including object placement, image blending, image harmonization, and shadow generation, to improve the realism and consistency of augmented images